Light in air is incident on a piece of crown glass at an angle of 35 degreee relative to the normal. What is the angle of refraction in the glass?



1
Expert's answer
2021-09-26T18:48:03-0400

Index of refraction of crown glass is 1.5

Snell's law

"n_1 sin \theta_1 = n_2 sin \theta_2 \\

1 \times sin 35 = 1.5 \times sin \theta_2 \\

0.5736 = 1.5 \times sin \theta_2 \\

sin \theta_2 = \frac{0.5736}{1.5} =0.3824 \\

\theta_2 = 22.482 \; degree"